A 7.2 mu C point charge is placed in an external u
Answer #1

given

Q = 7.2 micro C

E = 77000 N/c

let d is the distance where net electric field is zero.

Apply, E = k*Q/d^2

d^2 = E/(k*Q)

d = sqrt(E/(k*Q))

= sqrt(77000/(9*10^9*7.2*10^-6))

= 1.09 m